OpenStudy (dmezzullo):

Would it be D? Evaluate x − 2y when x = 3 and y = −2 7 −7 1 −1

OpenStudy (anonymous):

C

OpenStudy (anonymous):

by which I mean 7

OpenStudy (dmezzullo):

Oh i see now

Parth (parthkohli):

...

OpenStudy (dmezzullo):

Wait How WOuld It Be 7?

OpenStudy (anonymous):

x = 3 y = -2 3 - 2(-2) = 3 + 4 = 7

OpenStudy (shubhamsrg):

whenever you are stuck at any question, or have any problems in studies, or even life, I'd advice you to contact @DLS :) he has helped me through out my life :')

OpenStudy (anonymous):